OpenStudy (melissa_something):

log (x+4) = log x + log 4

OpenStudy (melissa_something):

I got up to \[x/4 +1 =x\]

OpenStudy (mathstudent55):

Can you combine the logs on the right side?

OpenStudy (melissa_something):

I already did. Everything condenses to \[\log (x+4 / 4x) =0\]

OpenStudy (mathstudent55):

\(\log (x + 4) = \log x + \log 4\) \(\log (x + 4) = \log 4x\) \(4x = x + 4\)

OpenStudy (melissa_something):

Oh wow, that looks easier...

OpenStudy (melissa_something):

What do I do after?

OpenStudy (mathstudent55):

Subtract x from both sides.

OpenStudy (mathstudent55):